iF A PERSON HAVING X PENS HE IS GOING TO ONE CLASS.IN THE CLASS 19 STUDENTS ARE THERE HE DISTRIBUTED PENS EQUALLY 3 PENS ARE REMAINING AFTER HE IS GOING TO ANOTHER CLASS IN THAT CLASS 23 STUDENTS ARE THERE HE DISTRIBUTED PENS EQUALLY THEN THERE ARE 5 PENS REMAINING FIND THE WHAT IS THE VALUE OF X?

Grade:11

1 Answers

Sher Mohammad IIT Delhi
10 years ago
x=19a+3
x=23b+5
19a-23b-2=0
a= (2+23b)/19, should be a integer,
b=9 and a=11 is one solution, x=19*11+3=212
the number of pens he have is 212
sher mohammad
B.Tech, IIT Delhi